Lakeland Chemical manufactures a product called Zing. Direct materials are added at the beginning of the process, and conversion activity occurs uniformly throughout production. The beginning work-in-process inventory is 80% complete with respect to conversion; the ending work-in-process inventory is 15% complete. The following data pertain to May: Units Work in process, May 1 36,500 Units started during May 70,500 Units completed and transferred out 77,500 Work in process, May 31 29,500 Total Direct Materials Conversion Costs Costs: Work in process, May 1 $ 50,750 $ 20,300 $ 30,450 Costs incurred during May 247,930 75,800 172,130 Totals $ 298,680 $ 96,100 $ 202,580 Using the weighted-average method of process costing, the cost per unit of direct materials is: Multiple Choice $0.90. None of the answers is correct. $1.00. $1.08. $1.20.
Answer #1

Equivalent unot fo material = 77500+29500 = 107000

Cost per equivalent unit of material = 96100/107000 = 0.90

So answer is a) $0.90